OpenStudy (superhelp101):

4. What is the possible discriminant of the graph? a parabola opening up with vertex near negative 1.2 comma negative 9.1. a y intercept of negative 6 -11 zero 25 73 5. What is the possible discriminant of the graph? a parabola opening up x intercepts near negative 7 and negative 3 -13 zero 15 16

OpenStudy (superhelp101):

@Abhisar do you get this?

OpenStudy (superhelp101):

@mathmate @Hero

OpenStudy (mathmate):

I'll take a look at that. It's possible to calculate, but only approximately, because the given information is "near"...

OpenStudy (superhelp101):

ok! I would appreciate it much! :)

OpenStudy (superhelp101):

@mathmate

OpenStudy (mathmate):

Sorry, I was busy with other ongoing replies. Will be working on it soon.

OpenStudy (superhelp101):

thx

OpenStudy (mathmate):

Finally back!

OpenStudy (mathmate):

Recall the quadratic formula: \(\huge x=\frac{-b\pm \sqrt{b^2-4ac}}{2a}\)

OpenStudy (mathmate):

If we consider \(\Delta=b^2-4ac \) then \(\huge x=\frac{-b\pm \sqrt{\Delta}}{2a}\) and consequently \(\huge x_2-x_1= \frac{\sqrt{\Delta}}{a}\) We can read off \(x_2-x_1\) from the graph. Can you figure out how you can find "a" to complete the solution?

OpenStudy (mathmate):

@superhelp101 AFK for a few minutes, brb.

OpenStudy (superhelp101):

I don't really understand

OpenStudy (mathmate):

Are you ok till the quadratic formula?

OpenStudy (superhelp101):

yes

OpenStudy (mathmate):

Subsequently, I have replaced the discriminant by \(\Delta\) to focus on the problem.

OpenStudy (mathmate):

ok up to that point?

OpenStudy (mathmate):

@superhelp101 still there?

OpenStudy (superhelp101):

yep

OpenStudy (superhelp101):

did you get 73 for the #4?

OpenStudy (mathmate):

I have a hard time reading the two zeroes, so cannot say. Can you tell me what the two zeroes are, or better still, what is x2-x1?

OpenStudy (superhelp101):

(-3.306,0) and (.806, 0) 4.112 is the difference

OpenStudy (superhelp101):

I not sure if that's exactly what you asked?

OpenStudy (mathmate):

Yes, 4.112 is what I need. I have a=2.15 approx. So according to the previous formula, \(x_2-x_1=\sqrt{\Delta}/a\) so \(\Delta = (a(x_2-x_1))^2=a^2(x_2-x_1)^2\) =78 approx.

OpenStudy (superhelp101):

so do you think 73 makes sense

OpenStudy (mathmate):

I rechecked a=2.1528, can you recheck (x2-x1). I don't think there should be that big a difference (78-73)/73=7%.

OpenStudy (superhelp101):

but I did .806-(-3.306) and got 4.112

OpenStudy (mathmate):

Thank you. I will do some rechecking and get back to you.

OpenStudy (superhelp101):

ok. will it take long thou lol?

OpenStudy (mathmate):

No, actually I just finished checking.

OpenStudy (superhelp101):

ok

OpenStudy (mathmate):

I'll try a different but hopefully better approach.

OpenStudy (mathmate):

If the roots are given to 3 decimal places, we'd better use them.

OpenStudy (mathmate):

The answer \(should\) be more accurate.

OpenStudy (mathmate):

We write f(x)=a(x-x1)(x-x2) Since f(0)=-6 (y-intercept), we get a=-6/f(0)=-6/(x1*x2)=2.252 approximately.

OpenStudy (mathmate):

ok so far? Please check my numbers.

OpenStudy (superhelp101):

yes just checked.

OpenStudy (mathmate):

So next we expand f(x) above numerically to get:

OpenStudy (mathmate):

\(f(x)=2.252x^2+5.629x-6\) from which we can easily calculate the discriminant \(b^2-4ac\).

OpenStudy (superhelp101):

31.685641-4(2.252)(-6)

OpenStudy (superhelp101):

31.685641-54.048

OpenStudy (mathmate):

Haha, this is even bigger. Since some of the numbers are approximate, perhaps the answers are based on some calculations of the approximate numbers. We had 78 before using those numbers. We cannot go wrong with what we've got, because they are accurate to 3 figures!

OpenStudy (mathmate):

The only approximate number we used was f(0)=y-intercept = -6.

OpenStudy (superhelp101):

what should we do now?

OpenStudy (mathmate):

I would take 73. If they say it's not correct, you can show your prof what we've done, and let him explain. The only catch is if your prof showed you a different method using the approximate values (h,k) and f(0), which \(might\) give a different answer.

OpenStudy (superhelp101):

ok thxx!!

OpenStudy (mathmate):

You're welcome! :)